Roll a fair die and let X be the square of the score that show up. Write down the
probability distribution of X hence compute
P(X <15)
and
P(3<= X< 30)

Answer by Boreal(15235)

the sample set is 1/4/9/16/25/36
So p(X<15) is 0.5
p(3<=x<30) is 2/3 , involving 4/9/16/25 but not 1 and 36.
p(1)=1/6=p(4)=p(9)=p(16)=p(25)=p(36)